# Mitigating Unsafe Feedback with Learning Constraints

### v2

@@ Conclusion
In this work, we find that LLMs are vulnerable to RPAs (though not as much as to HFTAs) even when a small percentage of preference labels in an RLHF training dataset are flipped, which is concerning given the popularity of feedback collection mechanisms, preferences datasets and the use of RLHF for training LLMs. Based on this finding, we develop a conceptual framework for adapting current defences against harmful fine-tuning attacks based on Constrained MDPs (CMDP). We unite various defences under a common framework of learning constraints both “online” or “explicit” and “offline” or “implicit”, and show how they fit in the CMDP framework. We explore the efficacy of these defences under a variety of settings, finding that several online defences are generally protective against RPAs without detracting from the ability to learn a harmless preference task. While offline defences provide some protection, they are generally not effective against RPAs pointing to the necessity of future research when the defence assumption does not allow the defender to apply an online defence.
We have shown a surprising result that some defences which do mitigate optimizing harmful reward end up performing “harmless reward hacking” and explained of why we might expect this to be the case by showing that optimal CMDPs and reward hacking have equivalent outcomes.
@@ Acknowledgments
We acknowledge the support of the Killam foundation, Vector Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI), Digital Research Alliance of Canada, and Research Nova Scotia.

### v3

@@ Conclusion
In this work, we find that LLMs are vulnerable to RPAs, which is concerning given the popularity of feedback collection mechanisms, preferences datasets, and the use of RLHF for training LLMs. Based on this finding, we develop a conceptual framework for adapting current defences against harmful fine-tuning attacks based on Constrained MDPs (CMDP). We unite various defences under a common framework of learning constraints both “online” or “explicit” and “offline” or “implicit”, and show how they fit in the CMDP framework. We explore the efficacy of these defences under a variety of settings, finding that several online defences are generally protective against RPAs without detracting from the ability to learn a harmless preference task. While offline defences provide some protection, they are generally not effective against RPAs pointing to the necessity of future research when the defence assumption does not allow the defender to apply an online defence.
We have shown a surprising result that some defences which do mitigate optimizing harmful reward end up performing “harmless reward hacking” and explained why we might expect this to be the case by showing that optimal CMDPs and reward hacking have equivalent outcomes.
@@ Acknowledgments
We thank Elahe Rahimi for reading an early version of this paper and providing valuable comments. We acknowledge the support of the Killam foundation, Vector Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI), Digital Research Alliance of Canada, and Research Nova Scotia.